OpenStudy (anonymous):

The half-life of ^60Co (Cobalt-60) is 5.27 years. How long does it take for 50 grams of ^60Co to decay to 20 grams?

OpenStudy (anonymous):

\[^{60}Co\]

OpenStudy (anonymous):

\[20=50\times (\frac{1}{2})^{\frac{t}{5.27}}\] solve for t

OpenStudy (anonymous):

\[\frac{2}{5}=.4=(\frac{1}{2})^{\frac{t}{5.27}}\] \[\frac{t}{5.27}=\frac{\ln(.4)}{\ln(.5)}\] \[t=\frac{5.27\ln(.4)}{\ln(.5)}\] then a calculator

OpenStudy (anonymous):

t = 1.3219, thanks so much! :)